Circular No. 8527

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V378 SERPENTIS R. W. Russell, D. K. Lynch, and R. J. Rudy, The Aerospace Corpo-

ration, obtained a 0.8–5.0-μm spectrogram of V378 Ser (cf. IAUC 8505, 8506) on Apr. 18.7 UT using the Infrared Telescope Facility (+ SpeX). The object showed a large number of symmetric, Gaussian-like emission lines (2700 km/s FWHM). H I dominated the spectrum with lines of the Pa, Br, Pf, and Hu series. He I was surprisingly weak, the 1.0830-μm line being about a third of the equivalent width of Paγ. No evidence was found of the CO 2.3- or 4.6-μm bands. Reddening determined from the O I lines was E(B − V) = 0.74. There was no indication of thermal emission from dust.

SUPERNOVAE 2005bi, 2005bu, AND 2005bz Further to IAUC 8523, K. Shimasaki and W. Li report the LOSS dis-

covery of an apparent supernova on unfiltered KAIT images taken on Apr. 29.51 (at mag 19.1) and May 12.44 UT (mag 19.0). SN 2005bz is located at α = 18^{h}13^{m}01^{s}.73, δ = +29^{o}41′51′′.5 (equinox 2000.0), which is 1′′.0 east and 4′′.4 north of the nucleus of UGC 11162. KAIT images taken on 2004 Aug. 19.31 (limiting mag 20.0) and 2005 Apr. 22.50 (limiting mag 19.0) showed nothing at this position.

A. V. Filippenko, R. J. Foley, D. Pooley, and R. Chornock, University of California, Berkeley, report that inspection of CCD spectra (range 320– 930 nm), obtained on May 11 UT with the Keck I 10-m telescope (+ LRIS), shows that SN 2005bi (IAUC 8511) is of type II, perhaps 1.5–2 months after the explosion. Lines of H I, Fe II, Ca II, and other species are present — in most or all cases with well-developed P-Cyg profiles. Spectra of SN 2005bu (IAUC 8519) show it to be of type Ia, roughly 3–4 weeks past maximum brightness.

COMET C/1996 N3 (SOHO) Another tailless SOHO comet (continuation to IAUC 8525):

Comet 1996 UT α_{2000} δ_{2000} Inst. F MPEC
C/1996 N3 July 3.383 6^{h}44.^{m}8 +22^{o}19′ C2/3 RK 2005-H24

K. Battams writes that C/1996 N3 was one of the brightest Meyer-group comets that he has seen (reaching mag 5.7 in C2 images on July 3.608 UT at 4.0R_{⊙}), and it didn’t fade noticeably until well past perihelion.

2005 May 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green